# A Framework of SO(3)-equivariant Non-linear Representation Learning and its Application to Electronic-Structure Hamiltonian Prediction

### v4

@@ I Discussion on Time Costs
J Empirical Study on Combining our Method with Approximately SO(3)-equivariance Framework
@@
While non-strict SO(3)-equivariance, which may limit the depth of theoretical exploration, is not the main focus of this study aiming at bridging rigorous SO(3)-equivariance with the non-linear expressive capabilities of neural networks, considering that they remain of interest in a few numerical computation applications where precision is highlighted over strict equivariance, we also conduct empirical study combining our method with approximately SO(3)-equivariant techniques. Taking the Monolayer MoS2 ($MM$) database as a case study, we evaluate the performance of combining our trace supervision and gradient induction method (TraceGrad) with the an approximately equivariant approach HarmoSE (Yin et al., 2024). We here take HarmoSE as the backbone encoder, and yields features by TraceGrad to enrich its representations. The experimental results in Table 8 and Fig. 7 demonstrate that TraceGrad significantly enhances the accuracy of HarmoSE, surpassing DeepH-2 and achieving SOTA results. Both DeepH-2 and HarmoSE sacrificed strict SO(3)-equivariance to fully release the expressive capabilities of graph Transformers, aiming for the ultimate in prediction accuracy. Despite this, our method still manages to significantly exceed their accuracy, further confirming the superiority and generality of our method in learning expressive representations of physical systems.
@@
K Future Work
In future research, various extensions are conceivable across theoretical, methodological, and application fields:
First, from a theoretical and methodological perspective, we can generalize SO(3) group symmetry to a broader range and more complex groups in deep learning research. This will enable deep learning frameworks to accommodate richer mathematical structures and physical quantities. Second, from the application perspective, the current methodology extends beyond predicting electronic-structure Hamiltonians and is capable of generally predicting multiple physical quantities and properties that are equivariant. Furthermore, as our approach offers a promising general solution to the critical dilemma between equivariance and non-linear expressiveness for the deep learning paradigm, the methodology is not confined to physical research alone and may find applications in areas such as robotics, autonomous vehicles, and motion tracking systems. For instance, in the vision tasks of autonomous vehicles (Wang et al., 2023), where there is a high demand for robustness against coordinate transformations of cameras, the mainstream approach has been to achieve approximate SO(3)-equivariance through data augmentation. However, this does not guarantee absolute reliability and safety. Our work holds great potential to construct deep models with high generalization performance under the premise of strict SO(3)-equivariance, potentially revolutionizing techniques in such tasks. Looking forward to future endeavors, the foundational and pivotal role of this work is clear.

### v5

@@ I Theoretical Analysis on Computational Complexity
L Empirical Study on Combining our Method with Approximately SO(3)-equivariance Framework
@@
While non-strict SO(3)-equivariance, which may limit the depth of theoretical exploration, is not the main focus of this study aiming at bridging rigorous SO(3)-equivariance with the non-linear expressive capabilities of neural networks, considering that they remain of interest in a few numerical computation applications where precision is highlighted over strict equivariance, we also conduct empirical study combining our method with approximately SO(3)-equivariant techniques. Taking the Monolayer MoS2 ($MM$) database as a case study, we evaluate the performance of combining our trace supervision and gradient induction method (TraceGrad) with the an approximately equivariant approach HarmoSE (Yin et al., 2024). We here take HarmoSE as the backbone encoder, and yields features by TraceGrad to enrich its representations. The experimental results in Table 10 and Fig. 7 demonstrate that TraceGrad significantly enhances the accuracy of HarmoSE, surpassing DeepH-2 (Wang et al., 2024b) and achieving SOTA results. Both DeepH-2 and HarmoSE sacrificed strict SO(3)-equivariance to fully release the expressive capabilities of graph Transformers, aiming for the ultimate in prediction accuracy. Despite this, our method still manages to significantly exceed their accuracy, further confirming the superiority of our method in learning expressive representations of physical systems for Hamiltonian prediction.